OpenStudy (anonymous):

Which of the following is equal to square root of the cube root of 2? 2 to the power of 2 over 3 2 to the power of 3 over 2 2 to the power of 1 over 3 2 to the power of 1 over 6

OpenStudy (anonymous):

Think of it this way One by one: cube root of 2 is 2^(1/3) Then square root it so it's (2^(1/3))^(1/2) = 2^(1/3)(1/2) = 2^(1/6)! therefore the answer is D

OpenStudy (anonymous):

thank you